# INTRABENCH: Interactive Radiological Benchmark

### v1

@@ 2.3 Human Effort Proxy
A model’s performance is highly dependent on the effort a human puts into initial prompting and refinement of the predicted masks. Generally, the effort required for 3D methods is less than that for 2D methods, although the strategies mentioned above significantly reduce the effort of 2D methods substantially. We aimed to establish a general measure of the effort a method would require from a human user. A more formalized mathematical approach involves assigning degrees of freedom (DoF) to each interaction. For instance, a point corresponds to 3 DoF, a 2D box has 5 DoF (requiring selection of the z-axis and two 2D points), and a 3D box consists of 6 DoF. However, point interpolation has 9 DoF, whereas point propagation only has 5 DoF, since it requires just the axial coordinate rather than both minimum and maximum points with 3 DoF each. From the user’s perspective, however, identifying the z-coordinate demands the same level of effort as selecting a 3D coordinate by clicking at the target structure’s endpoint along the z-axis. Similarly, an arbitrary scribble has significantly more DoF than a straight or parabolic line, yet the difference in effort for the user is minimal. Therefore, we define user effort in terms of the number of interactions required for a specific task. While not an exact measure, this method offers the most practical estimation of the actual effort involved from the user’s perspective.
2.4 Interactive methods
@@
In our comprehensive benchmark, we include various interactive segmentation methods. Fig. 1 illustrates the types of prompts each method supports. Iterative refinement is only possible for methods that allow a (previously predicted) mask as a prompt.
@@
SAM is the most prominent model from the natural image domain, that inspired many researchers to evaluate and adapt it to the domain of radiological medical images. It was trained on iteratively generated and curated 1B masks and 11M images, but not explicitly on radiological images. META’s Segment Anything Model was the first to popularize interactive segmentation models (Kirillov et al., 2023).
@@
SAM2 is an extension of SAM that was trained on even more images and introduced support for video data (Ravi et al., 2024).
@@
MedSAM is an adaptation of SAM that fine-tuned SAM’s weights on 1,570,263 image-mask pairs from the medical domain. It supports only a single forward pass without refinement and is limited to box prompts (Ma et al., 2024).
@@
SAM-Med 2D is another adaptation of SAM, fine-tuned on 4.6 million images with 19.7 million masks from the medical domain. Unlike MedSAM, it supports points, boxes, and mask prompts, allowing for refinement (Cheng et al., 2023).
@@
SAM-Med 3D incorporates a transformer-based 3D image encoder, 3D prompt encoder, and 3D mask decoder. It was trained from scratch using 22,000 3D images and 143,000 corresponding 3D masks and supports point and mask prompts and also allows for refinement (Wang et al., 2024).
@@
SAM-Med 3D Turbo is an updated version of SAM-Med 3D trained on a larger dataset collection of 44 datasets for improved performance. It supports the same prompt styles as SAM-Med 3D (Wang et al., 2024).
@@
SegVol is an interactive 3D segmentation model based on a 3D adaptation of a ViT (Dosovitskiy, 2020) that was trained on 96K unlabelled CT images and fine-tuned with 6K labeled CT images. It supports points and bounding boxes as spatial prompts but does not allow iterative refinement (Du et al., 2024).
@@
Aside from these models there exist other notable interactive models, such as Vista3D (He et al., 2024), 3D Sam Adapter (Gong et al., 2023) and Prism (Li et al., 2024). However, while being promptable, they are closed-set, i.e. not trained to segment any arbitrary prompted structure. Subsequently, they were not considered for this benchmark.
@@ 2.5 Datasets
Dataset selection was a non-trivial problem for this benchmark: While models that were originally introduced in the natural image domain rarely see any radiological 3D data, the medical counterparts were often trained on all publicly available datasets that the authors could obtain. For example, MedSAM was trained using more than 60 publicly available datasets (Ma et al., 2024). Although these methods conducted their final validation on excluded datasets or at least on separate test subsets of images, the test datasets vary between models. As a result, identifying annotated datasets with interesting target structures that were not part of any of the included methods’ training datasets has proven challenging.
Nevertheless, we assembled a diverse collection of ten lesser-known or recently released public datasets featuring various pathologies and organs, including CT and MRI image modalities. Specific details of these are provided in Table 1. To enhance reproducibility and eliminate barriers of entry for non-domain experts, we automated the dataset download and preprocessing, minimizing any required domain knowledge to use the benchmark. However, due to the sparsity of labeled datasets, we urge developers to exclude these datasets from their train dataset selection, as inclusion would compromise the integrity of a clean evaluation through IntRaBench.
@@
D1 MS Lesion (Muslim et al., 2022) MRI (T2 Flair) MS Lesions 60
D2 HanSeg (Podobnik et al., 2023) MR (T1) 30 Organs at risk 42
D3 HNTSRMFG (Wahid et al., 2024) MRI (T2) Oropharyngeal cancer & metastatic lymph nodes 135
D4 RiderLung (Zhao et al., 2015) CT Lung lesions 58
D5 LNQ (Dorent et al., 2024) CT Mediastinal lymph nodes 513
D6 LiverMets (Simpson et al., 2023) CT Liver metastases 171
D7 Adrenal ACC (Moawad et al., 2023) CT Adrenal tumors 53
D8 HCC Tace (Moawad et al., 2021) CT Liver, Liver tumors 65
D9 Pengwin (Liu et al., 2023) CT Bone fragments 100
D10 Segrap (Luo et al., 2023) CT 45 Organs at risk 30
@@ 2.6 Evaluation
All interactive segmentation methods identify their target structure based on a spatial prompt, inherently resulting in instance segmentation. As a result, we evaluate on an instance-by-instance basis. Unlike in object detection, each prompt already provides information on the localization of the target structure, making detection metrics like F1-Score irrelevant. Subsequently, we rely solely on the Dice Similarity Coefficient (DSC) score as a metric. The instance-wise DSC metric is then averaged per case (i.e. per image volume), and further aggregated across all cases in the dataset, as recommended by Maier-Hein et al. (2024). For better presentation, we averaged the DSC across all classes of a dataset and also specified how many human interactions are simulated.
3 Experiments
We evaluate all seven models across various initial prompting scenarios under realistic and unrealistic effort settings. Following this, we conduct interactive experiments to simulate human refinement of model predictions. Due to the vast amount of data, we only provide a condensed version of the results for easier insights. Detailed results and the number of human interactions are provided in the Appendix B.
@@ 3.1 Initial prediction
As an upper baseline, we begin with an idealized and unrealistic scenario where each slice is prompted individually for all 2D models. In this setting, we evaluate different numbers of point prompts per slice (PPS), as well as alternating positive and negative prompts ($\pm$ PPS), and slice-wise box prompts with varying numbers of boxes per slice (BPS). Figure 4 shows that models employing box prompts achieved significantly higher average Dice scores, with SAM2 demonstrating the strongest performance across all models. Conversely, point-based prompts performed poorly, particularly for small target regions, such as small MS lesions in dataset D1 (see Table 2). SAM Med2D outperforms non-medical models for point prompts. Although including positive and negative prompts and increasing the number of point prompts led to improvements, these were minor compared to the marked superiority of box-based prompts. These results highlight the limitations of point prompts, especially in cases involving small or complex anatomical structures, and emphasize the robustness of box prompts in achieving higher segmentation accuracy.
Figure 4: Unrealistic prompting of 2D Boxes each slice performs best. When comparing model’s prompted with one Box Prompts Per Slice (BPS) (left) with various Point Prompts per Slice (PPS) (right) boxes perform better. While alternating positive and negative points (dashed lines) is slightly superior to only positive points the gap between points and boxes remains large.
@@
To simulate a human-in-the-loop scenario, we evaluate various prompting strategies that avoid slice-by-slice interaction. As described in Section 2, for 2D models, we test point and box interpolation, as well as propagation, using different numbers of initial prompts. For 3D models, we explore varying numbers of Point prompts Per Volume (PPV) and 3D box prompts. Fig. 5 presents the following key findings:
@@
    For all models, box interpolation with 3 or 5 initial 2D boxes is sufficient to achieve results similar to slice-wise box prompting (BPS).
@@
    For SAMMed 2D, using 3 points with simple point interpolation achieves results comparable to prompting every slice.
@@ 3.2 Interactive Refinement
Finally, we evaluate the performance of the models during iterative refinement. For 2D models, this involves prompting on a slice-by-slice basis. As illustrated in Fig. 6 (left), adding refinement prompts to each slice results in a substantial performance boost. Although the proposed scribble-based refinement consistently improves outcomes, it does not achieve the same level of improvement as adding a prompt to every slice, which is expected since not all slices receive new prompts during the scribble refinements. We observed that for 2D models, it is crucial to provide the initial prompts again for each of the refinement steps. 2D models tend to over-segment the target, filling the entire slice foreground. The absence of the initial prompt leads to a complete loss of target location information, as the initial predicted mask is highly inaccurate. Our refinement likely generates negative additional prompts due to the large number of false positive pixels. In Table 5, we present refinement results from initial predictions produced by Box Interpolation. In this case, we did not include the previous point in the iterative prompts, which resulted in a performance decline during refinement.
For 3D models, iterative refinement also led to consistent performance improvements. Both randomly sampled prompts and those derived from refinement scribbles improved performance with each refinement iteration. Although SegVol initially performs best during initial prediction, it lacks support for further refinement. In contrast, SamMed 3D Turbo - worse in the initial prediction - surpasses SegVol, which was prompted using points, after several refinement steps.
Figure 6: All models demonstrate significant improvements from iterative refinement. Results for 2D models are shown on the left, and for 3D models on the right. Dashed lines represent the use of the proposed scribble refinement. While the unrealistic scenario of one refinement point per slice yields better performance, the proposed scribble refinement consistently enhances results across iterations for 2D models.
3.3 Discussion and Conclusion
In this paper we introduced IntRaBench  and with it, compared the performance of 2D and 3D interactive segmentation models in 3D medical imaging. We provide a holistic and transparent overview of the current state-of-the-art and highlight key findings that offer practical insights:
@@ 3.3 Discussion and Conclusion
    Bounding Boxes Outperform Points: Bounding boxes consistently outperform point-based inputs by providing better spatial context, which leads to improved segmentation accuracy, especially for complex structures in radiological images. Point-based prompts lack this context, resulting in poorer performance.
@@
    Iterative Refinement is Essential: The ability to iteratively refine segmentations significantly enhances model performance, particularly in challenging cases. Models that allow multiple rounds of corrections show better accuracy, making this feature crucial for clinical applications. For example SegVol reached highest performance in a static setting, however SamMed 3D Turbo is able to exceed SegVol given a few interactions, highlighting the importance of refinement.
@@
    Realistic 2D prompting can match unrealistic prompting: Our introduced realistic prompting styles are able to reach and match unrealistic prompting 2D prompting methods, see Fig. 5. This unlocks 2D methods for actual clinical workflows without any performance penalties.
@@
    Points fail for difficult and small structures: Contrary to claims in previous literature, point-based methods fail, likely due to previous work training and evaluating their methods on simpler target structures.
Implications
IntRaBench suggests that bounding boxes and iterative refinement should be prioritized in the design of segmentation models for medical imaging, particularly when addressing complex radiological images. Furthermore, it underscores the importance of including diverse, difficult tasks in training data to improve model generalization for clinical use. It is also crucial to test 2D models in scenarios that simulate real human interaction, ensuring that segmenting a volumetric image does not require unreasonable effort by prompting the model slice-by-slice.
@@
A key limitation of this work is that it only simulates ”Human Effort”. While this approach provides valuable insights into model performance by providing a proxy for the simulated ”Human Effort”, it falls short of capturing the full complexity and practical challenges of real clinical applications. As future work, a comprehensive study involving clinicians is essential to assess different prompting strategies in real-world environments. Such a study should not only evaluate segmentation performance but also measure the time required for annotation, offering critical insights into the practical feasibility and efficiency of these models in clinical practice.
To conclude, our proposed IntRaBench presents a powerful tool for the future of interactive segmentation research in medical imaging, serving as a catalyst for innovative solutions by enabling a fair and reproducible comparison between leading methods. One of the standout potentials is its ability to streamline the evaluation of both 2D and 3D segmentation models, allowing for more realistic and clinically relevant testing conditions. By focusing on human interaction and the efficiency of iterative refinement, IntRaBench opens new avenues for research, including understanding the impact of different interaction strategies and how they reduce clinician effort. Not only does this benchmark address the existing gaps in evaluation standardization, but it also offers a unique opportunity to refine segmentation performance on pathologies often overlooked, such as small lesions. The open-source nature of the benchmark further encourages continuous contributions, allowing researchers to test new methods and prompting strategies seamlessly within this framework. Future work using IntRaBench can reveal novel insights into the balance between performance and clinician involvement. This potential to improve real-world clinical applications, especially by reducing the labor intensity of medical professionals, marks IntRaBench as a crucial tool in catalyzing meaningful research progress.
Appendix A Model specifications
A.1 SAM

### v2

@@ 2.3 Human Effort Proxy
A model’s performance is highly dependent on the effort a human invests into the initial prompting and the refinement of the predicted masks. Generally, the effort of using 3D methods is lower than the effort of using 2D methods, although the strategies mentioned above significantly reduce the effort of 2D methods.
@@
We aimed to establish a quantitative measure of the effort a method would require from a human user, however, various issues arise: Considering a formalized mathematical approach where effort is quantified through the degrees of freedom (DoF) of each interaction. A point would correspond to 3 DoF, a 2D box 5 DoF (requiring selection of the z-axis and two 2D points), and a 3D box corresponds to 6 DoF. The point interpolation corresponds to 9 DoF, whereas point propagation would only be 5 DoF, since it requires just the axial coordinate rather than both minimum and maximum points with 3 DoF each. From the user’s perspective, however, identifying the z-coordinate demands the same level of effort as selecting a 3D coordinate by clicking at the target structure’s endpoint along the z-axis. Similarly, an arbitrary scribble has significantly more DoF than a straight or parabolic line, yet the difference in effort for the user is minimal. Subsequently, this approach was discarded.
Instead, we define user effort in terms of the number of interactions (i.e. clicks of a user) required for a specific prompt. This leads to e.g. a 2D and 3D click being equivalent in cost. While this is not an exact measure, it offers the most pragmatic estimation of the actual effort involved from a user’s perspective.
2.4 Interactive Segmentation methods
@@
In our benchmark, we include various interactive segmentation methods. Fig. 2 illustrates the types of prompts each method supports. Iterative refinement is only possible for methods that allow a (previously predicted) mask as a prompt.
@@
SAM is the most prominent model from the natural image domain, that inspired many researchers to evaluate and adapt it to the domain of radiological medical images. It was trained on iteratively generated and curated 1B masks and 11M images, but not explicitly on radiological images. It was the first to popularize interactive segmentation models [20].
@@
SAM2 is an extension of SAM that was trained on even more images and introduced support for video data [33].
@@
MedSAM is an adaptation of SAM that fine-tuned SAM’s weights on 1,570,263 image-mask pairs from the medical domain. It supports only a single forward pass without refinement and is limited to box prompts [24].
@@
SAM-Med 2D is another adaptation of SAM, fine-tuned on 4.6 million images with 19.7 million masks from the medical domain. Unlike MedSAM, it supports points, boxes, and mask prompts, allowing for refinement [3].
@@
SAM-Med 3D incorporates a transformer-based 3D image encoder, 3D prompt encoder, and 3D mask decoder. It was trained from scratch using 22,000 3D images and 143,000 corresponding 3D masks and supports point and mask prompts and also allows for refinement [39].
@@
SAM-Med 3D Turbo is an updated version of SAM-Med 3D trained on a larger dataset collection of 44 datasets for improved performance. It supports the same prompt styles as SAM-Med 3D [39].
@@
SegVol is an interactive 3D segmentation model based on a 3D adaptation of a ViT [7] that was trained on 96K unlabelled CT images and fine-tuned with 6K labeled CT images. It supports points and bounding boxes as spatial prompts but does not allow iterative refinement [9].
@@
Aside from these models there exist other notable interactive models, such as Vista3D [12], 3D Sam Adapter [11] and Prism [21]. However, while being promptable, they are closed-set, i.e. not trained to segment any arbitrary prompted structure. Subsequently, they were not considered for this benchmark.
@@ 2.5 Datasets
Dataset selection was a non-trivial problem for this benchmark: While models that were originally introduced in the natural image domain rarely see any radiological 3D data, the medical counterparts were often trained on all publicly available datasets that the authors could obtain. For example, MedSAM was trained using more than 60 publicly available datasets [24]. Although these methods conducted their final validation on excluded datasets or at least on separate test subsets of images, the test datasets vary between models. As a result, identifying annotated datasets with interesting target structures that were not part of any of the included methods’ training datasets has proven challenging.
Nevertheless, we assembled a diverse collection of ten lesser-known or recently released public datasets featuring various pathologies and organs, including CT and MRI image modalities. Specific details of these are provided in Tab. 1. While the quantity of cases may seem low, many contain multiple instances or target structures. Moreover, due to the pathological nature of many they represent the currently common size of datasets in the medical domain. To enhance reproducibility and eliminate barriers of entry for non-domain experts, we automated the dataset download and pre-processing, minimizing any required domain knowledge to use the benchmark. However, due to the sparsity of labeled datasets, we urge developers to exclude these datasets from their train dataset selection, as inclusion would compromise the integrity of a clean evaluation through RadioActive.
@@
D1 MS Lesion [31] MRI (T2 Flair) MS Lesions 60
D2 HanSeg [32] MR (T1) 30 Organs at risk 42
D3 HNTSRMFG [38] MRI (T2) Oropharyngeal cancer & metastatic lymph nodes 135
D4 RiderLung [44] CT Lung lesions 58
D5 LNQ [6] CT Mediastinal lymph nodes 513
D6 LiverMets [36] CT Liver metastases 171
D7 Adrenal ACC [29] CT Adrenal tumors 53
D8 HCC Tace [28] CT Liver, Liver tumors 65
D9 Pengwin [22] CT Bone fragments 100
D10 Segrap [23] CT 45 Organs at risk 30
@@ 2.6 Evaluation
All interactive segmentation methods identify their target structure based on a spatial prompt, inherently resulting in instance segmentation. As a result, we evaluate on an instance-by-instance basis. Unlike in object detection, each prompt already provides information on the localization of the target structure, making detection metrics like F1-Score irrelevant. Subsequently, we rely solely on the Dice Similarity Coefficient (DSC) score as a metric. The instance-wise DSC metric is then averaged per case (i.e. per image volume), and further aggregated across all cases in the dataset, as recommended by Maier-Hein et al. [25]. For better presentation, we averaged the DSC across all classes of a dataset and also specified how many human interactions were simulated.
3 Experiments and Results
In our experiments, we evaluate all seven models across all datasets. First, we evaluate the performance during the initial prompting step, and following this, we evaluate the iterative refinement performance of methods that support this. For each of these, we test models in both, realistic and high-effort prompting schemes. To enhance readability, we present the results together with the primary findings in the main manuscript, while additional, more granular insights are deferred to Sec. 7 in the appendix.
@@ 3.1 Initial prediction
As an upper baseline, we begin with an idealized and unrealistic scenario where each slice is prompted individually for all 2D models. In this setting, we evaluate different numbers of point prompts per slice (PPS), as well as alternating positive and negative prompts ($\pm$ PPS), and slice-wise box prompts with varying numbers of boxes per slice (BPS). Figure 5 shows that models employing box prompts achieved significantly higher average Dice scores, with SAM2 demonstrating the strongest performance across all models. Conversely, point-based prompts performed poorly, particularly for small target regions, such as small MS lesions in dataset D1 (see Table 2 in the appendix). Although including positive and negative prompts and increasing the number of point prompts led to improvements, these were minor compared to the marked superiority of box-based prompts. Interestingly, generalist models not trained with a particular focus on medical data performed worse with point prompts per slice but outperformed models trained on medical data when using box prompts. This may be because models trained with medical data are better at resolving the ambiguity of point prompts, especially given the often low contrast between pathologies and surrounding tissue. In contrast, SAM and SAM2 models excel in object identification when using box prompts.
Figure 5: Unrealistic prompting of 2D Boxes each slice performs best. 2D models prompted with one Box Prompt Per Slice (BPS), indicated with a star, outperform models prompted with various Point Prompts per Slice (PPS, line plots). Drawing a 2D box accounts for 2 interactions similar to providing 2 points. While alternating positive and negative points (dashed lines) is slightly superior to only positive points the gap between points and boxes remains large.
@@
To simulate a human-in-the-loop scenario, we evaluate various prompting strategies that avoid slice-by-slice interaction. As described in Sec. 2, for 2D models, we test point and box interpolation, as well as propagation, using different numbers of initial prompts. For 3D models, we explore varying numbers of Point prompts Per Volume (PPV) and 3D box prompts. Fig. 6 presents the following key findings:
@@
    For all models, simple prompt interpolation methods achieve similar performance as the corresponding prompt per-slice methods. For example for SAM-Med 2D, 3 Point Interpolation even outperforms prompting 1 PPS having a mean DSC of 26.08 compared to a DSC of 23.13. This is because we select central points for the interpolation methods instead of random foreground points. Another example is SAM prompted using 5 Box Interpolation, that has a mean DSC of 56.9 compared to a DSC of 57.84 when prompted with a box for each slice.
@@ 3.2 Interactive Refinement
Finally, we evaluate the performance of the models during iterative refinement. For 2D models, this involves predicting on a slice-by-slice basis. As illustrated in Fig. 7 (top), adding refinement prompts to each slice results in a substantial performance boost. While the proposed scribble-based refinement consistently improves outcomes, it does not achieve the same level of improvement as adding a prompt to every slice, which is expected since not all imperfect slices are guaranteed to receive new prompts during the scribble refinements. Moreover, we observed that for 2D models, it is crucial to provide the initial prompts again for each of the refinement steps. In the appendix in Table 5, we present results without including the previous point in the iterative refinement step, which resulted in a performance decline during refinement. 2D models tend to over-segment the target, filling the entire slice foreground. The absence of the initial prompt leads to a complete loss of target location information, as the initial predicted mask is highly inaccurate. Our refinement likely generates negative additional prompts due to the large number of false positive pixels. Qualitative examples of this failure case are presented in the appendix in Sec. 7 in the appendix.
For 3D models, iterative refinement also led to consistent performance improvements. Both randomly sampled prompts and those derived from refinement scribbles improved performance with each refinement iteration. Although SegVol initially performs best for the initial prediction and achieves a mean DSC of 43.32 prompted with 5 centered point prompts, it lacks support for further refinement. In contrast, SAM-Med 3D Turbo - worse in the initial prediction with a Dice of 40.89 - surpasses SegVol after several refinement steps reaching a DSC of 47.4.
Figure 7: All models demonstrate significant improvements from iterative refinement. Dashed lines represent the use of the proposed scribble refinement. While the unrealistic scenario of one refinement point per slice yields better performance, the proposed scribble refinement consistently enhances results across iterations for 2D models.
4 Discussion and Conclusion
In this paper we introduced RadioActive  and with it, compared the performance of 2D and 3D interactive segmentation models in 3D medical imaging. We provide a holistic and transparent overview of the current state-of-the-art and highlight key findings that offer practical insights:
@@ 4 Discussion and Conclusion
    Realistic 2D prompting can match unrealistic prompting: Our introduced realistic prompting styles can match unrealistic 2D prompting methods, see Fig. 6. This unlocks 2D methods for actual 3D medical image workflows without any major performance penalties.
@@
    SAM 2 outperforms all medical 2D & 3D models: When prompting SAM2 through a realistic prompting scheme such as the Box Interpolation Sam 2 is superior to all medical 2D & 3D models. It even surpasses models after unrealistic iterative refinements.
@@
    Iterative Refinement is essential: The ability to iteratively refine segmentations significantly enhances model performance, particularly in challenging cases. Models that allow multiple rounds of corrections show better accuracy, making this feature crucial for clinical applications. For example, SegVol reached the highest performance in a static setting for point prompts, however, SAM-Med 3D Turbo is able to exceed SegVol given a few interactions, highlighting the importance of refinement.
@@
    Points fail for difficult and small structures. Contrary to claims in previous literature [3, 34], point-based methods fail, likely due to previous work training and evaluating their methods on simpler target structures [24, 3, 34].
  5. 5.
    Bounding boxes outperform points. Consistent with previous studies [34, 24, 9, 27], we find that Bounding boxes consistently outperform point-based inputs by providing better spatial context, which leads to improved segmentation accuracy, especially for complex structures in radiological images. Point-based prompts lack this context, resulting in poorer performance.
Limitations
@@
While the RadioActive benchmark addresses many issues, our work has limitations. Currently, we only approximate Human Effort. While the pragmatic approach provides valuable insights into model performance in relation to effort, it falls short of capturing the full complexity and practical challenges of real clinical applications. As future work, a comprehensive study involving clinicians is essential to assess the efficacy of different prompting strategies in real-world environments. Such a study should not only evaluate segmentation performance but also measure the time required for annotation, offering critical insights into the practical feasibility and efficiency of these models in clinical practice. Currently, the benchmark includes ten datasets, encompassing a variety of pathologies and selected anatomical structures in MRI and CT images. We acknowledge that the quantity of images in these datasets is relatively small—a common limitation for most rare lesion datasets—all large, however, most larger publicly available datasets have already been utilized in training the methods making this challenging. Another limitation of this work is that, given the rapid development of new interactive segmentation methods, very recent approaches, such as Wong et al. [41], are not yet included. However, we consider this a strength, as it highlights the need for a fair and open benchmark to keep up with ongoing advancements [26]. The published codebase is designed to seamlessly integrate new prompting strategies and models, enabling rapid adaptation and extension.
Conclusion
The open-source nature of RadioActive  serves as a catalyst for interactive segmentation research in 3D medical imaging by enabling a fair and reproducible comparison between existing methods. Future work can build upon and extend this framework to investigate crucial questions in interactive segmentation. Through this, we believe RadioActive will contribute to improving real-world clinical application of interactive segmentation methods, facilitating a reduction of labor for medical professionals and accelerating meaningful clinical research.
Supplementary Material
5 Model specifications
5.1 SAM